# Numerical tiling-based simulations of decoherence in multifield models of inflation

### v2

@@ Abstract
In previous work, we developed a method for computing two-point correlators by decomposing the mode degrees of freedom into fast and slow components. Building on this framework, we present a numerical implementation to study the evolution of primordial scalar perturbations under controlled state deformations induced by the simplest environment corrections from the Lindblad equation. Our approach generalizes to an arbitrary number of degrees of freedom and does not rely on the slow-roll approximation. The computational routine is numerically efficient and allows users to configure arbitrary sequences of decoherence events, with full control over their duration, shape, amplitude and effective wavelength range. The resulting outputs are compatible with nonlinear numerical codes, enabling studies of how decoherence effects propagate during reheating.
@@ I Introduction
Building on this separation method, we modify the evolution scheme to solve the simplest realization of the Lindblad equation presented in [18, 19, 20, 21, 22], and extend it to the multifield case. The numerical routines are designed to accommodate an arbitrary sequence of decoherence events (or “accidents”) acting on different $k$-modes at different times, without impacting on computational performance. This sequence defines the time-dependent power spectrum of the environmental operators and determines the resulting deformed state of each mode. The duration, amplitude and range of wavelengths of these events can be tuned to generate distinctive features in the evolution of Gaussian states and in the resulting power spectrum of primordial curvature fluctuations.
@@
The remainder of this paper is organized as follows. In Section II, we review the open quantum system formalism in the context of inflationary models driven by a single scalar field. We also introduce the foundations of our fast–slow scale separation scheme and compute the curvature power spectrum under the influence of multiple accidents introducing decoherence. In Section III, we extend our component separation method to incorporate decoherence effects in multifield inflation models. Our results demonstrate that it is possible to generate features in the power spectrum of primordial curvature fluctuations across any range of wavelengths. Using cross-correlations, we illustrate how decoherence propagates between isocurvature and isentropic degrees of freedom. In Appendix A, we show that our decomposition scheme can also be applied to other forms of coloring transformations beyond the Cholesky decomposition. In Appendix B, we construct three-dimensional field realizations from the outputs of our numerical routine. These field realizations can serve as initial conditions for simulations of the reheating phase. Finally, in Section IV, we discuss potential future applications of these routines and present our conclusions.
@@ II.2 Two-point correlators of a system coupled to an environment
where $\alpha_{\Gamma}(\ln\ell_{j},N_{i})$ is a dimensionless function specifying the strength of the environmental effect for the tile $(i,j)$. This expression is obtained from the amplitude prescription for a single accident given, for example, in [35], after transforming to the $(N,\ln\ell)$ variables. Here, $\ln\ell_{j}$ denotes the logarithm of the physical length scale at which the accident is most effective, $N_{i}$ is the number of e-folds where the tile is located, and $p$ controls the steepness of the event. A key feature of our approach is that $\alpha_{\Gamma}(\ln\ell_{j},N_{i})$ may take positive and negative values, representing, respectively an increase or a decrease in the number of available states encoded in each Wigner ellipse. For the evolution of Gaussian states described in Eqns. (29a29c), negative values $\alpha_{\Gamma}$ do not violate the positivity of the Wigner function, except when the covariance matrix becomes singular. The functions $T(N)$ and $\kappa(\ln\ell)$ behave as window functions centered at $N_{i}$ and $\ln\ell_{j}$, respectively, with the widths of the corresponding length and time bands for the tile $(i,j)$, given by $\Delta N_{ij}$ and $\Delta\ell_{ij}$.
@@ II.3 Implementation of fast and slow component separation for single field models
$\frac{\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle}{\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}}=\frac{L_{k}|_{\tau_{0}}}{L^{4}_{k}}\left[1+\Delta_{p}(k,\tau)\right]\,$, (43)
@@
$\bigg[\frac{\rm d^{2}}{\rm d\tau^{2}}+\left(k^{2}-\frac{z^{\prime\prime}}{z}\right)-\frac{L_{k}|_{\tau_{0}}}{L^{4}_{k}}\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}$ (44)
$-\frac{L_{k}|_{\tau_{0}}}{L^{4}_{k}}\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}\Delta_{p}(k,\tau)\bigg]L_{k}=0\,.$
@@
$\bigg[\frac{\rm d^{2}}{\rm d\tau^{2}}+\left(k^{2}-\frac{z_{\rm eff}^{\prime\prime}}{z_{\rm eff}}\right)-\frac{L_{k}|_{\tau_{0}}}{L^{4}_{k}}\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}\bigg]L_{k}=0\,$,
$\frac{z_{\rm eff}^{\prime\prime}}{z_{\rm eff}}\equiv\frac{z^{\prime\prime}}{z}+\frac{L_{k}|_{\tau_{0}}}{L^{4}_{k}}\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}\Delta_{p}(k,\tau)\,.$ (45)
@@
which are only different from the standard definitions by a rescaling factor of $k$. Using the evolution of field correlators at fixed wavenumbers, Figure 5, panels (c1) and (c2), illustrate how the squeezing parameters evolve away from their accident-free values as the $k$-modes pass through the accident/anti-accident pairs, and continue evolving up to the end of inflation. [^4] More explicitly, deviations are measured with respect to the squeezing parameters $\varphi^{(0)}_{k}$ and $r^{(0)}_{k}$, which arise from computing the evolution of correlators in the accident-free scenario, defined by $\mathcal{F}(k,\tau)=0$. From both subpanels, it is clear that the most significant effects occur when the modes traverse each individual (anti-) accident within the pair. As an ancillary file attached to this paper, we include an animation showing the evolution of the Wigner ellipse and its projections over the normalized axes $k^{-1/2}\pi_{k}$ and $k^{1/2}v_{k}$, as an arbitrary $k$-mode traverses a decoherence event to produce a squeezed state. While arbitrary arrays of decoherence events can be used to generate squeezed states, they are more effective at inducing variations in the area of the Wigner ellipse than at producing strong squeezing.
@@
In the right panel of Figure 6, we depict the deformations of the primordial power spectrum caused by the sunflower configuration of tiles displayed in the left panel. As in Figure 4, panel (b), our results show that it is possible to produce features that may either enhance or suppress power in any band of wavenumbers. This behavior is consistent with the effective field theory approach to inflation and is particularly relevant for mimicking effects of radiatively stable models of inflation that exhibit different forms of instabilities (e.g., [40, 41, 42, 16, 32]). It should be emphasized that, although some of the power spectra depicted in this work are not consistent with current observational constraints, their role is to exemplify the flexibility of the framework in accommodating a wide variety of feature profiles at different wavelengths.
The results shown in Figures 5 and 6 suggest that decoherence accidents can be configured to reconstruct the primordial power spectra of more intricate inflationary models. In Figure 7, we construct a sequence of decoherence events designed to reproduce the characteristic features of a multifield inflationary model mode by mode. The power spectrum of such a model exhibits multiple deformations induced by instabilities sourced by the field-space curvature. As an example, we produced the spectrum of a two-field model where the background dynamics is governed by the potential
@@
The right panel compares the curvature spectrum obtained from this deformation procedure with both the original and target spectra. The results shown in this figure exploit an expected degeneracy: initial-state deformations in a simple inflationary scenario and dynamical features of a more intricate model can, in principle, produce (nearly) indistinguishable primordial power spectra. Although the resulting spectrum exhibits small but non-negligible deviations from the multifield spectrum, this indicates that more precise convergence techniques – such as those based on machine learning algorithms – may be required to reduce these discrepancies to machine precision. Nevertheless, our results are sufficient to demonstrate that, starting from a basic inflationary model, this formalism can reproduce features in the power spectrum characteristic of more sophisticated models using relatively simple sequences of decoherence events.
@@ III.3 Implementation of a dynamical coloring scheme via Cholesky decomposition
Some care must be taken when applying this implementation, since the original set of transport equations in Eqns. (82a82c) still forms a third-order system with respect to $\left\langle\hat{\mathbf{\Sigma}}_{\rm S}^{(vv)}\right\rangle$, which – when written as in Eqns. (78b) and (84) – cannot be order-reduced. At super-horizon scales, this system exhibits a sign flip in the diagonal elements of $\bm{\mathcal{Z}}$ – which are positive, by definition – leading to numerical instabilities that manifest particularly in scenarios where the background trajectories undergo sharp turns. Despite this, the framework retains enough flexibility to generate arbitrarily large features in the primordial power spectrum, provided that (a) the model potentials remain sufficiently smooth and (b) decoherence and recoherence events occur at subhorizon scales (as developed in section II for one field). Moreover, since part of the purpose of this approach is to mimic Gaussian state deformations of more complicated dynamical setups using smooth, well-behaved models as a starting point, these conundrums do not pose severe limitations to achieving this objective. Extending this approach to accommodate potentials that induce more intricate deformations of the background evolution will be the subject of future work.
@@ III.4 Results: evolution of the Wigner ellipse and generation of features in the power spectrum
In addition to the deformation of initial states to reproduce the primordial spectrum of more complex (i.e., with more fields, intricate potentials, nontrivial effects from field geometry or combinations of all of these) inflationary models, the inclusion of decoherence within the multifield framework raises a fundamental question: how do these field effects manifest and propagate across coupled degrees of freedom? Addressing this question requires characterizing the dynamics of each Wigner ellipse as it undergoes the simplest sequences of decoherence events. To this end, we adopt the same parameterization introduced in section II (Eqns. (33) and (53)) for the single-field case, which specifies the amplitude, location, duration, and wavelength band over which each event is effective.
@@
Figure 11: Mode correlators resulting from the implementation of the two decoherence channels built from the environmental operators $\hat{I}^{\rm S}_{\parallel}$ and $\hat{I}^{\rm S}_{\perp}$ defined in Eqns. (109a109b). Panel (a): Mode-injection scheme with an array of decoherence events, using the same tiling in the $(N,\ln\ell)$ plane for both operators. The highlighted mode (purple dashed line) shows the evolution of mode correlators across each decoherence region. Panels (b1) and (b2): Parallel $(\mathcal{P}_{\mathcal{RR}})$ and perpendicular $(\mathcal{P}_{\mathcal{SS}})$ spectra for the $\hat{I}^{\rm S}_{\parallel}$ and $\hat{I}^{\rm S}_{\perp}$ operators. Each operator imprints features primarily on its corresponding spectrum, with residual structures in $\mathcal{P}_{\mathcal{RR}}$ arising from small correlations between isocurvature and adiabatic modes. Panel (c): Evolution of the $\mathcal{P}_{\cal RR}$ $k$-mode highlighted in Panel (a), as it crosses parallel and perpendicular decoherence events. The inset shows the variation $\Delta\mathcal{P}_{\cal RR}$ relative to the accident-free scenario $(\hat{I}^{\rm S}=0)$, indicating that $\mathcal{P}_{\cal RR}$ modes grow when crossing the $\hat{I}_{\perp}^{\rm S}$ configuration, which is consistent with the presence of the small-amplitude features visible (in green) in panel (b2). Panel (d): Cross correlators $(\mathcal{P}_{\mathcal{RS}})$ for the same setup. The large spike (in blue) marks a sign flip rather than a change in correlation magnitude. The inset illustrates how increasing the field coupling enhances the amplitude of cross correlations. Features are present in all of the cases.
@@
An animation accompanying our results in Figure 9, provided as an ancillary file, illustrates the full evolution of the joint Wigner function as the modes traverse decoherence/recoherence events for the same tile parameters considered in this figure. In this animation, the evolution also includes the deformation of the ellipse during horizon crossing. A second animation shows the generation of squeezed states under a different sequence of tiles for the single-field version of this model (i.e., $V(\phi)=\lambda\phi^{4}/4$).
@@
Considering the interaction operator definitions introduced above, our results in Figure 11 display the power, cross-correlations and the dynamical effects from decoherence events in the evolution of the mode correlators defined in Eqns. (110a110b) for arbitrary values of $k$. These results are obtained by configuring a fixed sequence of decoherence events sourced separately by the parallel and perpendicular operators. Individual tiles are built following the parallelogram parameterization introduced in Eq. (53).

### v3

@@ Abstract
In previous work, we developed a method for computing two-point correlators by decomposing the mode degrees of freedom into fast and slow components. Building on this framework, we present a numerical implementation to study the evolution of primordial scalar perturbations under controlled state deformations induced by the simplest environment corrections from the Lindblad equation. The primary contribution of this work is the development of a numerically stable and flexible framework for implementing open-system effects in inflationary perturbations, rather than the extraction of concrete predictions for the primordial spectrum. Our approach generalizes to an arbitrary number of degrees of freedom and does not rely on the slow-roll approximation. The computational routine is numerically efficient and allows users to configure arbitrary sequences of decoherence events, with full control over their duration, shape, amplitude and effective wavelength range. The resulting outputs are compatible with nonlinear numerical codes, enabling studies of how decoherence effects propagate during reheating.
@@ I Introduction
Building on this separation method, we modify the evolution scheme to solve the simplest realization of the Lindblad equation presented in [18, 19, 20, 21, 22], and extend it to the multifield case. The numerical routines are designed to accommodate an arbitrary sequence of decoherence events (or “accidents”) acting on different $k$-modes at different times, without impacting on computational performance. This sequence defines the time-dependent power spectrum of the environmental operators and determines the resulting deformed state of each mode. The duration, amplitude and range of wavelengths of these events can be tuned to generate distinctive features in the evolution of Gaussian states and in the resulting power spectrum of primordial curvature fluctuations. This framework is agnostic about the origins of decoherence, and the physical interpretation of specific arrays of decoherence events is deferred for future work.
@@
The remainder of this paper is organized as follows. In Section II, we review the open quantum system formalism in the context of inflationary models driven by a single scalar field. We also introduce the foundations of our fast–slow scale separation scheme and compute the curvature power spectrum under the influence of multiple accidents introducing decoherence. In Section III, we extend our component separation method to incorporate decoherence effects in multifield inflation models. Our results demonstrate that it is possible to generate features in the power spectrum of primordial curvature fluctuations across any range of wavelengths. Using cross-correlations, we illustrate how decoherence propagates between isocurvature and isentropic degrees of freedom. Following Eq. (104), this section also includes a detailed discussion of the limitations and regime of validity of the present implementation. In Appendix A, we show that our decomposition scheme can also be applied to other forms of coloring transformations beyond the Cholesky decomposition. In Appendix B, we construct three-dimensional field realizations from the outputs of our numerical routine. These field realizations can serve as initial conditions for simulations of the reheating phase. Finally, in Section IV, we discuss potential future applications of these routines and present our conclusions.
@@ II.2 Two-point correlators of a system coupled to an environment
where $\alpha_{\Gamma}(\ln\ell_{j},N_{i})$ is a dimensionless function specifying the strength of the environmental effect for the tile $(i,j)$. This expression is obtained from the amplitude prescription for a single accident given, for example, in [35], after transforming to the $(N,\ln\ell)$ variables. Here, $\ln\ell_{j}$ denotes the logarithm of the physical length scale at which the accident is most effective, $N_{i}$ is the number of e-folds where the tile is located, and $p$ controls the steepness of the event. It is important to emphasize that the tiling formalism is not intended to model a specific microscopic environment, but rather to span a space of effective open-system histories consistent with Gaussian evolution. Nevertheless, tile arrays can be used to mimic concrete physical situations. For example, sharp background trajectory turns [38, 32] may be represented by transient tiles, while heavy-field thresholds [39] in natural or chaotic models can be replicated by localized tile arrays. In this sense, the tiling approach should be viewed as a diagnostic tool rather than a first-principles model of environmental interactions during inflation. A key feature of our approach is that $\alpha_{\Gamma}(\ln\ell_{j},N_{i})$ may take positive and negative values, representing, respectively an increase or a decrease in the number of available states encoded in each Wigner ellipse. For the evolution of Gaussian states described in Eqns. (29a29c), negative values $\alpha_{\Gamma}$ do not violate the positivity of the Wigner function, except when the covariance matrix becomes singular. The functions $T(N)$ and $\kappa(\ln\ell)$ behave as window functions centered at $N_{i}$ and $\ln\ell_{j}$, respectively, with the widths of the corresponding length and time bands for the tile $(i,j)$, given by $\Delta N_{ij}$ and $\Delta\ell_{ij}$.
@@ II.3 Implementation of fast and slow component separation for single field models
$\frac{\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle}{\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}}=\frac{L^{4}_{k}|_{\tau_{0}}}{L^{4}_{k}}\left[1+\Delta_{p}(k,\tau)\right]\,$, (43)
@@
$\bigg[\frac{\rm d^{2}}{\rm d\tau^{2}}+\left(k^{2}-\frac{z^{\prime\prime}}{z}\right)-\frac{L^{4}_{k}|_{\tau_{0}}}{L^{4}_{k}}\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}$ (44)
$-\frac{L^{4}_{k}|_{\tau_{0}}}{L^{4}_{k}}\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}\Delta_{p}(k,\tau)\bigg]L_{k}=0\,.$
@@
$\bigg[\frac{\rm d^{2}}{\rm d\tau^{2}}+\left(k^{2}-\frac{z_{\rm eff}^{\prime\prime}}{z_{\rm eff}}\right)-\frac{L^{4}_{k}|_{\tau_{0}}}{L^{4}_{k}}\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}\bigg]L_{k}=0\,$,
$\frac{z_{\rm eff}^{\prime\prime}}{z_{\rm eff}}\equiv\frac{z^{\prime\prime}}{z}+\frac{L^{4}_{k}|_{\tau_{0}}}{L^{4}_{k}}\langle\tilde{\chi}^{\prime}_{\rm S},\tilde{\chi}^{\prime}_{\rm S}\rangle|_{\tau_{0}}\Delta_{p}(k,\tau)\,.$ (45)
@@
which are only different from the standard definitions by a rescaling factor of $k$. Using the evolution of field correlators at fixed wavenumbers, Figure 5, panels (c1) and (c2), illustrate how the squeezing parameters evolve away from their accident-free values as the $k$-modes pass through the accident/anti-accident pairs, and continue evolving up to the end of inflation. [^4] More explicitly, deviations are measured with respect to the squeezing parameters $\varphi^{(0)}_{k}$ and $r^{(0)}_{k}$, which arise from computing the evolution of correlators in the accident-free scenario, defined by $\mathcal{F}(k,\tau)=0$. From both subpanels, it is clear that the most significant effects occur when the modes traverse each individual (anti-) accident within the pair. As an ancillary file attached to this paper [42], we include an animation showing the evolution of the Wigner ellipse and its projections over the normalized axes $k^{-1/2}\pi_{k}$ and $k^{1/2}v_{k}$, as an arbitrary $k$-mode traverses a decoherence event to produce a squeezed state. While arbitrary arrays of decoherence events can be used to generate squeezed states, they are more effective at inducing variations in the area of the Wigner ellipse than at producing strong squeezing.
@@
In the right panel of Figure 6, we depict the deformations of the primordial power spectrum caused by the sunflower configuration of tiles displayed in the left panel. As in Figure 4, panel (b), our results show that it is possible to produce features that may either enhance or suppress power in any band of wavenumbers. This behavior is consistent with the effective field theory approach to inflation and is particularly relevant for mimicking effects of radiatively stable models of inflation that exhibit different forms of features (e.g., [43, 38, 44, 16, 32]). It should be emphasized that, although some of the power spectra depicted in this work are not consistent with current observational constraints, their role is to exemplify the flexibility of the framework in accommodating a wide variety of feature profiles at different wavelengths.
The results shown in Figures 5 and 6 suggest that decoherence accidents can be configured to reconstruct the primordial power spectra of more intricate inflationary models. As an example of this, in Figure 7, to illustrate the diagnostic power of this formalism, we construct a sequence of decoherence events designed to reproduce the characteristic features of a multifield inflationary model mode by mode. The power spectrum of such a model exhibits multiple deformations induced by instabilities sourced by the field-space curvature. As an example, we produced the spectrum of a two-field model where the background dynamics is governed by the potential
@@
The right panel compares the curvature spectrum obtained from this deformation procedure with both the original and target spectra. The results shown in this figure illustrate degeneracies: initial-state deformations in a simple inflationary scenario and dynamical features of a more intricate model can, in principle, produce (nearly) indistinguishable primordial power spectra. Although the resulting spectrum exhibits small but non-negligible deviations from the multifield spectrum, this indicates that more precise convergence techniques – such as those based on machine learning algorithms – may be required to reduce these discrepancies to machine precision. Rather than providing a new explanation for observed features, these results underscore the difficulty of uniquely attributing spectral signatures to background dynamics in the absence of independent constraints on environmental couplings, and highlight the limitations of inference based solely on the power spectrum.
@@ III.3 Implementation of a dynamical coloring scheme via Cholesky decomposition
Before presenting the numerical results, it is important to clarify the regime of validity of the transport system and the numerical limitations that arise in certain dynamical situations. Some care must be taken when applying this implementation, since the original set of transport equations in Eqns. (82a82c) still forms a third-order system with respect to $\left\langle\hat{\mathbf{\Sigma}}_{\rm S}^{(vv)}\right\rangle$, which – when written as in Eqns. (78b) and (84) – cannot be order-reduced. At super-horizon scales, this system exhibits a sign flip in the diagonal elements of $\bm{\mathcal{Z}}$ – which are positive, by definition – leading to numerical instabilities that manifest particularly in scenarios where the background trajectories undergo sharp turns. Despite this, the framework retains enough flexibility to generate arbitrarily large features in the primordial power spectrum, provided that (a) the model potentials remain sufficiently smooth and (b) decoherence and recoherence events occur at subhorizon scales (as developed in section II for one field). Moreover, since part of the purpose of this approach is to mimic Gaussian state deformations of more complicated dynamical setups using smooth, well-behaved models as a starting point, these implementation-related challenges do not pose severe limitations to achieving this objective. Extending this approach to accommodate potentials that induce more intricate deformations of the background evolution will be the subject of future work.
@@
Although the implementation presented here is restricted to trivial field-space metrics, the formalism can be extended to nontrivial field-space geometries without altering its essential structure. At the level of linear perturbations, we may introduce a local vielbein $(\Lambda_{A}^{i})$ on field space and perform a variable redefinition that maps the system to an approximately flat field-space frame, following the standard treatment (see e.g., [47]). Capital Latin indices label directions in the curved field space, while lowercase Latin indices correspond to the locally flat frame. Choosing the vielbein to satisfy parallel-transport equations along the background trajectory
$\mathcal{D}_{t}\Lambda^{i}_{A}\equiv\dot{\Lambda^{i}_{A}}+\Gamma^{C}_{AB}\dot{\phi}^{B}\Lambda^{i}_{C}=0\,$, (108)
allows for canonical quantization of the perturbation modes in this frame. This requires considering the vielbein as additional dynamical degrees of freedom. In this representation, the effects of field-space curvature are retained through corrections to the effective frequency matrix, while the kinetic structure remains canonical. Since the coloring transformation and tiling-based implementation of decoherence act directly on the mode evolution equations, their construction is not significantly modified in this frame. Sequences of decoherence events can therefore be implemented as in the previous sections, and the inverse vielbein may be used to reconstruct the perturbations in the original curved field-space basis.
@@ III.4 Results: evolution of the Wigner ellipse and generation of features in the power spectrum
In addition to the deformation of initial states to mimic the primordial spectrum of more complex (i.e., with more fields, intricate potentials, nontrivial effects from field geometry or combinations of all of these) inflationary models, the inclusion of decoherence within the multifield framework raises a fundamental question: how do these field effects manifest and propagate across coupled degrees of freedom? Addressing this question requires characterizing the dynamics of each Wigner ellipse as it undergoes the simplest sequences of decoherence events. To this end, we adopt the same parameterization introduced in section II (Eqns. (33) and (53)) for the single-field case, which specifies the amplitude, location, duration, and wavelength band over which each event is effective.
@@
Figure 11: Mode correlators resulting from the implementation of the two decoherence channels built from the environmental operators $\hat{I}^{\rm S}_{\parallel}$ and $\hat{I}^{\rm S}_{\perp}$ defined in Eqns. (110a110b). Panel (a): Mode-injection scheme with an array of decoherence events, using the same tiling in the $(N,\ln\ell)$ plane for both operators. The highlighted mode (purple dashed line) shows the evolution of mode correlators across each decoherence region. Panels (b1) and (b2): Parallel $(\mathcal{P}_{\mathcal{RR}})$ and perpendicular $(\mathcal{P}_{\mathcal{SS}})$ spectra for the $\hat{I}^{\rm S}_{\parallel}$ and $\hat{I}^{\rm S}_{\perp}$ operators. Each operator imprints features primarily on its corresponding spectrum, with residual structures in $\mathcal{P}_{\mathcal{RR}}$ arising from small correlations between isocurvature and adiabatic modes. Panel (c): Evolution of the $\mathcal{P}_{\cal RR}$ $k$-mode highlighted in Panel (a), as it crosses parallel and perpendicular decoherence events. The inset shows the variation $\Delta\mathcal{P}_{\cal RR}$ relative to the accident-free scenario $(\hat{I}^{\rm S}=0)$, indicating that $\mathcal{P}_{\cal RR}$ modes grow when crossing the $\hat{I}_{\perp}^{\rm S}$ configuration, which is consistent with the presence of the small-amplitude features visible (in green) in panel (b2). Panel (d): Cross correlators $(\mathcal{P}_{\mathcal{RS}})$ for the same setup. The large spike (in blue) marks a sign flip rather than a change in correlation magnitude. The inset illustrates how increasing the field coupling enhances the amplitude of cross correlations. Features are present in all of the cases.
@@
An animation accompanying our results in Figure 9, provided as an ancillary file [48], illustrates the full evolution of the joint Wigner function as the modes traverse decoherence/recoherence events for the same tile parameters considered in this figure. In this animation, the evolution also includes the deformation of the ellipse during horizon crossing. A second animation shows the generation of squeezed states under a different sequence of tiles for the single-field version of this model (i.e., $V(\phi)=\lambda\phi^{4}/4$).
@@
Considering the interaction operator definitions introduced above, our results in Figure 11 display the power, cross-correlations and the dynamical effects from decoherence events in the evolution of the mode correlators defined in Eqns. (111a111b) for arbitrary values of $k$. These results are obtained by configuring a fixed sequence of decoherence events sourced separately by the parallel and perpendicular operators. Individual tiles are built following the parallelogram parameterization introduced in Eq. (53).